# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v2

@@ 4.2 Results and Takeaways
Synthetic data. Figure 3 shows the evaluation bias – i.e., the proximity to true errorc – for four baseline metrics (Harrell’s CI, Uno’s CI, IPCW-integrated Brier Score, and MAE-Margin) and their dependent variants as a function of Kendall’s $\tau$. Assuming a known ground-truth copula (Archimedean family and dependency strength), we estimate $\hat{S}(t)$ using a CoxPH model to predict $N$ ISDs, where the risk scores are the negative predicted survival times. Overall, our dependent metrics show lower bias – i.e., closer proximity to the true error – when dependency is notable ($\tau>0.25$). For the ranking metrics (Harrell’s CI, Uno’s CI), biases are similar under low dependency ($\tau<0.25$) but improve significantly under high dependency ($\tau>0.5$). For accuracy metrics (IPCW-integrated Brier Score, MAE-Margin), our dependent variants consistently provide better error estimates regardless of dependency or copula family.
As a practical example, consider a cancer study, with high dependent censoring ($\tau=0.75$) from a Clayton copula between event time $E$ (cancer relapse) and censoring time ${C}$ (cancer death). Based on Figure 3, if the true CI of a CoxPH model is 0.65, then, if we evaluate the model using Harrell’s CI, we anticipate the CI to be 0.68 ($\pm$ 0.03). If we use CI-Dep, we anticipate the CI to be 0.67 ($\pm$ 0.02).
Semi-synthetic data. Table 1 presents the results in 3 semi-synthetic datasets: METABRIC, MIMIC-IV and SEER. The complete results for all 6 datasets are in Appendix G.1. In the small-sized METABRIC dataset, Harrell’s gives superior estimates of the true concordance index, and there is a performance degradation in using the dependent variant. We see that CI-Dep degrades the performance compared to Uno’s CI, in this case, while IBS-Dep improves on the IPCW version, and MAE-Margin and MAE-Dep give similar results. In the medium-sized MIMIC-IV dataset, Harrell’s CI is again better, but our IBS-Dep and MAE-Dep variants improve on the IBS-IPCW and margin MAE, respectively, in all cases, however, the pseudo MAE is the best (i.e., it has the lowest error). In the large-sized SEER dataset, there are few differences in discriminative performances for the three metrics, but our IBS-Dep improves on IPCW, while our MAE-Dep improves on pseudo and margin. More often than not, Harrell’s and Uno’s CI (IPCW) give better estimates than our CI-Dep in these datasets, while our IBS-Dep and MAE-Dep give notable improvements to their independent counterparts. See Appendix G.1 for additional results.
Dataset Strategy Discrimination (CI) Accuracy (IBS) Accuracy (MAE)
Harrell’s Uno’s Dependent IPCW Dependent Uncens. Hinge Pseudo Margin Dependent
METABRIC ($N$=1,102, $C$=3.8%) Original 0.10$\pm$0.08 0.39$\pm$0.14 0.48$\pm$0.25 1.23$\pm$0.85 1.17$\pm$0.85 3.24$\pm$0.55 1.40$\pm$0.21 0.25$\pm$0.21 0.31$\pm$0.23 0.33$\pm$0.22
Top 5 0.11$\pm$0.08 0.51$\pm$0.17 0.58$\pm$0.23 1.22$\pm$0.85 1.17$\pm$0.85 3.21$\pm$0.55 1.40$\pm$0.21 0.25$\pm$0.20 0.31$\pm$0.23 0.30$\pm$0.23
Top 10 0.10$\pm$0.08 0.43$\pm$0.11 0.49$\pm$0.16 1.23$\pm$0.85 1.17$\pm$0.85 3.26$\pm$0.54 1.40$\pm$0.21 0.25$\pm$0.21 0.32$\pm$0.23 0.31$\pm$0.23
Rand. 25% 0.10$\pm$0.08 2.34$\pm$0.68 2.31$\pm$0.70 1.25$\pm$0.88 1.18$\pm$0.88 3.45$\pm$0.48 1.40$\pm$0.21 0.27$\pm$0.22 0.34$\pm$0.23 0.33$\pm$0.23
MIMIC-IV (hospital) ($N$=6,780, $C$=74.3%) Original 0.74$\pm$0.43 1.27$\pm$0.83 2.10$\pm$1.54 8.75$\pm$1.50 6.57$\pm$1.44 3.48$\pm$0.25 6.44$\pm$0.27 3.11$\pm$0.27 4.54$\pm$0.26 4.36$\pm$0.24
Top 5 0.79$\pm$0.55 1.16$\pm$0.56 3.36$\pm$2.74 8.80$\pm$1.51 6.42$\pm$1.45 3.51$\pm$0.24 6.45$\pm$0.26 3.10$\pm$0.26 4.65$\pm$0.25 4.33$\pm$0.18
Top 10 0.71$\pm$0.47 1.10$\pm$0.65 2.46$\pm$1.69 8.73$\pm$1.50 6.57$\pm$1.44 3.50$\pm$0.27 6.45$\pm$0.27 3.11$\pm$0.27 4.60$\pm$0.28 4.40$\pm$0.25
Rand. 25% 1.44$\pm$0.58 1.77$\pm$1.02 2.70$\pm$1.44 10.38$\pm$0.76 7.91$\pm$0.90 3.57$\pm$0.20 6.51$\pm$0.19 3.06$\pm$0.19 4.79$\pm$0.19 4.62$\pm$0.20
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.77$\pm$0.10 0.57$\pm$0.14 0.76$\pm$0.13 0.33$\pm$0.15 0.25$\pm$0.11 2.59$\pm$0.07 2.50$\pm$0.04 2.47$\pm$0.07 1.48$\pm$0.06 1.44$\pm$0.05
Top 5 1.86$\pm$0.12 1.70$\pm$0.15 1.76$\pm$0.13 0.28$\pm$0.14 0.47$\pm$0.21 2.80$\pm$0.05 2.69$\pm$0.06 2.25$\pm$0.07 1.25$\pm$0.06 0.85$\pm$0.47
Top 10 0.81$\pm$0.08 0.90$\pm$0.09 1.10$\pm$0.07 0.33$\pm$0.15 0.25$\pm$0.11 2.61$\pm$0.08 2.51$\pm$0.04 2.46$\pm$0.06 1.47$\pm$0.05 1.43$\pm$0.05
Rand. 25% 0.89$\pm$0.52 1.80$\pm$1.23 1.68$\pm$1.21 0.45$\pm$0.21 0.35$\pm$0.17 2.75$\pm$0.11 2.49$\pm$0.08 2.50$\pm$0.14 1.49$\pm$0.09 1.47$\pm$0.11
@@
Table 1: Comparison of established survival metrics and the proposed dependent metrics in terms of the mean ($\pm$ SD.) bias in the semi-synthetic test sets, averaged over 5 learners and 5 experiments. CI and IBS results are multiplied by 100.
In general, a (censored) metric should be representative of the true metric, i.e., it should tell us which model makes the best survival predictions. Therefore, we create a new experiment as follows:
  • 1.
    In our semi-synthetic datasets, we are aware of the ground-truth (the actual event times) of all instances. For each model, we use this information to calculate the True CI, True IBS and True MAE results – that is, how good are the survival predictions using each metric, if all the instances were observed (not censored).
  • 2.
    For each true metric, we then rank the top-3 survival models – e.g., CoxPH, RSF and DeepSurv – by their true metric. In this example, CoxPH has the highest True CI, then RSF and then DeepSurv.
  • 3.
    For each censored metric (proposed and existing), we check if the metric gives the same ordering of survival models as its corresponding true metric – e.g., if True CI indicates that CoxPH, RSF and DeepSurv are the three best models, does CI-Dep indicate the same? This experiment is repeated 5 times with different data splits.
Figure 4 shows the number of times the resp. metric correctly (green dot) or incorrectly (red dot) identified the ordering of the top-3 survival learners in the METABRIC test set. See Appendix G for results in all semi-synthetic datasets.
Figure 4: Plot the ranking accuracy, i.e., the number of times the resp. metric correctly (green dot) or incorrectly (red dot) identified the true ordering of the top-3 survival learners, out of 5 experiments in the METABRIC test sets.
@@ 5 Discussion
How do the Metrics Work? As we do not have a formal proof of identifiability – i.e., we do not know the true relationship between $E$ and $C$ – we therefore study this question empirically by approximating $C_{\theta}$ and making a comparison between conventional metrics and our dependent variants. While the advantages to adopting our dependent metrics are evident in synthetic data, where we know the ground-truth copula and dependency level, these advantages vary with the dataset and its underlying distribution in the semi-synthetic datasets. Based on our observations, the most consistent trend is seen in the proposed MAE-Dep –- which is directly comparable to the MAE-Margin. Specifically, the MAE-Dep always outperforms its independent counterpart when censoring exceeds 20%. In these scenarios, MAE-Dep yields lower mean biases across five repeated experiments on all four dataset variants with over 20% censoring. This improvement is encouraging, as it holds across datasets with possibly varying characteristics and censoring mechanisms. However, when the censoring rate falls between 0% and 20%, the trend reverses slightly: the MAE-Dep tends to be marginally higher than the MAE-Margin.
What are the Limitations? Our metrics rely on a known or approximate copula to adjust for dependent censoring, but estimating the copula is inherently difficult due to the non-identifiability of competing-risks data [Tsiatis, 1975]. This poses a key limitation for applying our method to real-world datasets. However, there exist several strategies to approximate the copula using observed data: Emura and Chen [2016] proposed to divide the $N$ instances into $K$ groups of approximately equal sample sizes, and for each instance $i\in K$, define a prognostic index, $\text{PI}_{i}(C_{\theta},\bm{x}_{i},W)$, which predicts the outcome for $i$, where $W$ are the model parameters. Using cross-validation, we pick the copula $\hat{\theta}$ which maximizes the CI over possible values of Kendall’s $\tau$. Parikh et al. [2022] and Mahajan et al. [2024] proposed generating counterfactual data with characteristics similar to the observed data and evaluate the model performance on the simulated data using counterfactual metrics (such as the Survival-$\ell_{1}$ by Foomani et al. [2023]). Foomani et al. [2023] proposed to first get an intuition about the underlying copula (e.g., Clayton, Frank) by plotting the event and censoring distributions to visually inspect the curves and then go on to approximate $C_{\theta}$ using MLE.

### v3

@@ 4.2 Results and takeaways
Dependent metrics under increasing dependency. Figure 2 shows the evaluation bias – i.e., the absolute difference between the ground-truth metric and the censored metric – for four baseline metrics (Harrell’s CI, Uno’s CI, IPCW-integrated Brier Score, and MAE-Margin) and our dependent variants as a function of Kendall’s $\tau$. Assuming a known ground-truth copula, we estimate $\hat{S}_{T}(t|\bm{x})$ using a CoxPH model to make survival predictions, where the risk scores are the negative predicted times. Overall, our dependent metrics show lower bias – i.e., closer proximity to the true error – when dependency is notably ($\tau>0.25$) than their independent variants. For the ranking metrics (Harrell’s CI, Uno’s CI), biases are similar under low dependency ($\tau<0.25$) but improve significantly under high dependency ($\tau>0.5$). For the accuracy metrics (IPCW-integrated Brier Score, MAE-Margin), our dependent variants give equal or slightly better error estimates regardless of the dependency.
Dependent metrics under increasing censoring. Figure 3 shows the evaluation bias as a function of the censoring rate (0-80%). Here, we set a fixed Kendall’s $\tau$ to 0.5, draw 10,000 samples from the DGP over 100 experiments, and group the results by the censoring rates in the whole dataset. We assume a known ground-truth copula and estimate $\hat{S}_{T}(t|\bm{x})$ using a CoxPH model. Our dependent metrics are generally aligned with their independent counterparts under low censoring (0-20%), but consistently exhibit lower biases and less variance when censoring increases ($\geq 20\%$).
Figure 2: Plot of mean biases ($\pm$ 2 SD.) for baseline and proposed metrics as a function of the dependency under a known copula, averaged over 100 experiments ($N=10,000$, $d=10$).
Figure 3: Plot of mean biases ($\pm$ 2 SD.) for baseline and proposed metrics as a function of the censoring rate under a known copula, averaged over 100 experiments (Kendall’s $\tau$ = 0.5, $N=10,000$, $d=10$) and two copula models (Clayton and Frank).
5 Semi-synthetic experiments
5.1 Experimental setup
To evaluate our metrics on real datasets, we require knowledge of the true event times, which is not feasible due to censoring. To address this, we construct realistic semi-synthetic datasets where the covariates, event times, and censoring distribution reflect those of real-world datasets. Inspired by [23], we estimate a censoring distribution $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$ from a real dataset $\mathcal{D}$, remove censored instances to obtain $\mathcal{D^{\prime}}$, and then synthetically censor instances in $\mathcal{D^{\prime}}$ based on whether a sampled censoring time is earlier than the event time ($\tilde{c}_{i}<e_{i}$). We induce dependent censoring by omitting features in the original dataset based on one of three feature selection strategies. This ultimately gives us $\mathcal{D}^{\prime\prime}$, which has a synthetic but realistic censoring distribution, a restricted feature set, and known ground-truth event times, enabling us to compare the proposed and existing (censored) metrics to the true metrics. Figure 4 illustrates the evaluation pipeline. The remainder of this section outlines the practical details of our approach.
Figure 4: Evaluation pipeline. (1) Based on some raw, survival dataset $\mathcal{D}$, (2) we form $\tilde{\mathcal{D}}$ by flipping the event bit of $\mathcal{D}$ and then estimate the feature-dependent censoring distribution, $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$. (3) We then select a subset of features from $\mathcal{D}$ using some strategy (e.g., Top-K features), (4) and generate a ground-truth dataset, $\mathcal{D}^{\prime}$, by removing the censored instances from $\mathcal{D}$. (5) Putting all this together, we have a semi-synthetic dataset $\mathcal{D}^{\prime\prime}$, which is based on the ground-truth dataset, the learned censoring distribution, and the selected features. (6) We then train two Archimedean copula models to estimate $\hat{C}_{\theta}$ and (7) five survival learners to make predictions based on $\mathcal{D}^{\prime\prime}$. Finally, (8) we compare the baseline and proposed censored metrics (using $\hat{C}_{\theta}$) with the true evaluation metric.
Datasets (Step 1 in Figure 4): We apply the evaluation pipeline to 6 large, real-world datasets: METABRIC [35], 2 from MIMIC-IV (all causes and hospital death) [36], and 3 from SEER (death from brain, liver and stomach cancer) [37]. They differ in their number of instances ($N$), number of features ($d$) and censoring rates ($C$). See Appendix E.1 for more details.
Feature selection (Step 3 in Figure 4): To induce dependent censoring, we apply one of three feature selection strategies to $\mathcal{D}$: (1) Original, where all features are retained; (2) Top-$k$, where we select the $k$ most predictive features based on permutation feature importance [38] evaluated using Harrell’s CI [20]; and (3) Random-%, where a random subset of features is selected.
Copula model (Step 6 in Figure 4): We estimate $\hat{C}_{\theta}$ by training Clayton and Frank copulas with nonlinear Weibull proportional-hazards marginals for $T$ and $C$ on $\mathcal{D}^{\prime\prime}$. Each marginal uses a multilayer perceptron (MLP) with one hidden layer and ReLU activation function. Marginals and copulas are trained jointly via MLE (see Appendix E.3). The copula with the lowest validation NLL is selected for evaluation.
Survival learners (Step 7 in Figure 4): We compare the prediction results using 5 survival learners produced by the CoxPH [9], Gradient Boosting Survival Analysis (GBSA) [39], Random Survival Forests (RSF) [15], DeepSurv [13] and Multi-Task Logistic Regression (MTLR) [16]. These learners are trained to predict an individual survival distribution (ISD) for each instance in the test set. See Appendix E.2 for further details, including implementation details and hyperparameters.
We repeat the procedure outlined in Figure 4 over 5 experiments. In each of 5 experiments, we follow [34] and split $\mathcal{D}^{\prime\prime}$ into stratified training (70%), validation (10%), and test (20%) sets (seeds 0-4). After splitting, we impute missing values using the sample mean for real-valued covariates or the mode for categorical covariates based on the training set only. We then normalize numerical values and encode categorical features using a one-hot encoding strategy.
5.2 Results and takeaways
Due to space constraints, the main text presents only partial results for datasets with different characteristics, e.g., number of samples and censoring rates. See Appendix F.2 for the complete results, where we report the approximated copula $\hat{C_{\theta}}$ and also evaluate the ranking accuracy.
Discrimination. Table 1 presents the discrimiation (CI) results in terms of the evaluation bias in 3 semi-synthetic datasets: METABRIC, MIMIC-IV (all), and SEER (liver). In the METABRIC dataset, CI-Dep performs very similarly to Uno’s CI on Original, Top 5, and Top 10 strategies (bias $\approx$ 0.005) and slightly higher than Harrell’s CI. In the MIMIC-IV dataset, all the metrics are generally aligned with similar mean biases and error margins. In the SEER dataset, CI-Dep slightly diverges from CI-Uno, giving marginally higher bias on Top 10 and lower on Rand. 25% strategies. Overall, we see that CI-Dep does not significantly improve or degrade performances over the baseline metrics.
Accuracy. Table 2 presents the accuracy (IBS and MAE) results in terms of the evaluation bias the same datasets. In the METABRIC dataset, all IBS results are identical or near-identical across metrics and strategies and the MAE-Dep has slightly lower bias than Margin, which is its independent counterpart. In the MIMIC-IV dataset, IBS-Dep performs on par with KM, and MAE-Dep again closely matches other metrics, except Hinge. In the SEER dataset, IBS-Dep equals KM and IPCW, but the MAE-Dep performs better than Hinge, Pseudo and Margin in Top 5 and Top 10 strategies. We hypothesize that the similar biases are likely due to a lack of dependent censoring or the relatively low censoring rates (0–20%). In our synthetic experiments, more pronounced differences between the independent and dependent metrics typically emerged beyond 25% censoring (see Figure 3).
@@ 5.2 Results and takeaways
Table 1: Comparison of baseline and proposed discrimination metrics (CI) in terms of the mean ($\pm$ 2 SD.) bias, i.e., the proximity to the true error, averaged over 5 survival learners and 5 experiments.
Dataset Strategy Harrell Uno Dep.
METABRIC ($N$=1,102, $C$=3.8%) Original 0.001$\pm$0.002 0.004$\pm$0.003 0.005$\pm$0.005
Top 5 0.001$\pm$0.002 0.005$\pm$0.003 0.006$\pm$0.005
Top 10 0.001$\pm$0.002 0.004$\pm$0.002 0.005$\pm$0.003
Rand. 25% 0.001$\pm$0.002 0.023$\pm$0.014 0.023$\pm$0.014
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original 0.006$\pm$0.004 0.007$\pm$0.004 0.007$\pm$0.004
Top 5 0.011$\pm$0.004 0.016$\pm$0.005 0.016$\pm$0.005
Top 10 0.010$\pm$0.003 0.013$\pm$0.004 0.013$\pm$0.004
Rand. 25% 0.006$\pm$0.002 0.006$\pm$0.003 0.006$\pm$0.003
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.008$\pm$0.002 0.006$\pm$0.003 0.008$\pm$0.003
Top 5 0.019$\pm$0.002 0.017$\pm$0.003 0.017$\pm$0.003
Top 10 0.008$\pm$0.002 0.009$\pm$0.002 0.011$\pm$0.001
Rand. 25% 0.011$\pm$0.014 0.019$\pm$0.017 0.018$\pm$0.017
Table 2: Comparison of baseline and proposed accuracy metrics (IBS and MAE) in terms of mean ($\pm$ 2 SD.) bias, i.e., the proximity to the true error, averaged over 5 survival learners and 5 experiments. KM denotes the proposed IBS metric under the independent copula for comparison. Bold indicates best performing independent/dependent variant for the MAE.
IBS MAE
Dataset Strategy IPCW KM Dep. Hinge Pseudo Margin Dep.
METABRIC ($N$=1,102, $C$=3.8%) Original 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.246$\pm$0.414 0.314$\pm$0.459 0.328$\pm$0.431
Top 5 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.249$\pm$0.398 0.308$\pm$0.460 0.298$\pm$0.464
Top 10 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.248$\pm$0.418 0.317$\pm$0.462 0.308$\pm$0.465
Rand. 25% 0.012$\pm$0.018 0.012$\pm$0.018 0.012$\pm$0.018 1.397$\pm$0.416 0.266$\pm$0.435 0.342$\pm$0.463 0.328$\pm$0.467
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original 0.008$\pm$0.003 0.009$\pm$0.003 0.009$\pm$0.003 62.244$\pm$10.539 37.156$\pm$10.255 37.156$\pm$10.255 37.301$\pm$10.245
Top 5 0.008$\pm$0.004 0.011$\pm$0.004 0.011$\pm$0.004 65.626$\pm$11.943 39.558$\pm$11.741 39.558$\pm$11.741 39.717$\pm$11.743
Top 10 0.008$\pm$0.004 0.010$\pm$0.004 0.010$\pm$0.004 63.716$\pm$11.380 38.131$\pm$11.190 38.131$\pm$11.190 38.312$\pm$11.205
Rand. 25% 0.008$\pm$0.003 0.011$\pm$0.004 0.011$\pm$0.004 62.749$\pm$8.117 37.204$\pm$8.102 37.204$\pm$8.102 37.364$\pm$8.062
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.003$\pm$0.003 0.003$\pm$0.002 0.003$\pm$0.002 2.498$\pm$0.075 2.475$\pm$0.134 1.478$\pm$0.110 1.520$\pm$0.107
Top 5 0.003$\pm$0.003 0.003$\pm$0.002 0.003$\pm$0.002 2.688$\pm$0.122 2.246$\pm$0.134 1.250$\pm$0.123 1.226$\pm$0.219
Top 10 0.003$\pm$0.003 0.002$\pm$0.002 0.002$\pm$0.002 2.513$\pm$0.072 2.462$\pm$0.119 1.466$\pm$0.102 1.450$\pm$0.170
Rand. 25% 0.005$\pm$0.005 0.004$\pm$0.004 0.004$\pm$0.004 2.505$\pm$0.136 2.499$\pm$0.272 1.490$\pm$0.193 1.524$\pm$0.227
@@ 6 Discussion
Limitations. This work has limitations: First, our metrics rely on a known or well-approximated Archimedean copula (see the following paragraph) to adjust for dependent censoring. However, it is typically impossible to recover the true underlying copula due to the nature of competing-risks data: for each instance, we only observe one outcome (either event or censoring), but not both [40]. This is a key limitation of applying our metrics and of copula-based methods in general, since the ground-truth dependence structure cannot be directly observed or learned. Second, our metrics may be misspecified under a temporal distribution shift in the dependence structure. That is, if $P(T\mid X)$, $P(C\mid X)$, and $P(T,C\mid X)$ change over time, then the dependence structure evolves, and we would have to re-estimate the copula parameter. Third, successfully maximizing the likelihood (Equation 2.2) depends on the dataset characteristics – e.g., number of instances, censoring rate, etc. – which means that model estimates may have high variance in small sample sizes.
User guide. Our metrics require users to specify a dependence structure under an Archimedean copula. However, since no nonparametric test can conclusively detect the dependency between event and censoring, we instead offer a practical approach by estimating a likely dependence structure based on the log-likelihood: the question is not whether this will yield the optimal (latent) structure, but a structure better than simply assuming independence between event and censoring. The log-likelihood is a widely used loss function in survival analysis and is also a proper scoring rule [41]. We first establish a baseline using the independent copula model, then assess whether a dependent copula provides a better fit by approximating the data’s dependence structure. As a useful heuristic, $\mathcal{D}$ should contain 25% censoring or more, based on our observations, for dependent censoring to be relevant.
  1. 1.
    Given $\mathcal{D}$, split the data into stratified training, validation, and test sets. Define marginals (e.g., Weibulls) for event and censoring and maximize the dependent-likelihood (Equation 2.2) under a Clayton independent-copula $C_{\theta}$ based on the training data. This can be done using gradient descent, setting $\theta$ to a small positive constant (e.g.$10^{-4}$ such that $\theta>0$) and freezing its weight in the optimization procedure (see Appendix E.3). Record the validation loss and halt training when it stops improving. This will yield survival marginals, a fixed copula $\hat{C_{\theta}}\approx 0$ under independence, and some validation loss $\mathcal{L}_{\text{Indep}}$, which is our baseline.
  2. 2.
    Reset the marginal model weights and reinitialize $C_{\theta}$ as an Archimedean copula, but unfreeze its weight in the optimization procedure and maximize the dependent-likelihood (Equation 2.2) based on the training data. Again, record the validation loss and halt training when it stops improving. This will again yield survival marginals, a copula $\hat{C_{\theta}}$ under dependence, and some validation loss $\mathcal{L}_{\text{Dep}}$. Repeat Step 2 for each copula model to be evaluated.
  3. 3.
    Now, compare $\mathcal{L}_{\text{Indep}}$ and $\mathcal{L}_{\text{Dep}}^{k}$ for each of the $K$ trained copula models under dependence. If $\mathcal{L}_{\text{Dep}}^{k}<\mathcal{L}_{\text{Indep}}$, then the $k$-dependent model better captures the joint distribution, implying a better model fit. Alternatively, use a penalized likelihood criterion (e.g., AIC, BIC) and choose the model with the lower score to balance goodness-of-fit and complexity. Finally, use the approximated $\hat{C_{\theta}}$ from the chosen model to calculate our metrics. Recall that our metrics reduce to their independent variants when $C_{\theta}\approx 0$.
Broader impacts. Benchmarking survival models is inherently difficult due to censoring, as we do not have complete information about the survival times for some instances. Making wrong assumptions about the censoring mechanism can give a false impression of how well a model actually performs. This work advances the reliable evaluation of survival models under dependent censoring, which is important for real-world applications – such as biomedical and epidemiological research – where accurate predictions inform clinical decisions and public health policy.